# DuSEGO: Dual Second-order Equivariant Graph Ordinary Differential Equation

### v1

@@ 6.3. Molecular Data
Finally, we evaluate DuSEGO with the state-of-art baseline methods in molecular data, where we aim to predict molecular chemical properties on 3D molecular graphs.
@@
The QM9 dataset is widely used in computational chemistry, which is specifically designed for the development and testing of quantum chemistry methods. It consists of approximately 134,000 stable small organic molecules made up of C, H, O, N, and F. Each molecule in the dataset contains up to nine non-hydrogen atoms. The dataset includes a diverse set of molecular properties that have been pre-computed using high-level quantum mechanical theories. We imported the QM9 dataset partitions from (anderson2019cormorant, ) and sample 100K molecules for training, 18K for validation and 13K for testing (brandstetter2021geometric, ). Additionally all properties have been normalized by substracting the mean and dividing by the Mean Absolute Deviation (satorras2021n, ).
@@
Various widely used methods for molecule property prediction are deployed for comparison, such as DimeNet++ (gasteiger_dimenetpp_2020, ), SphereNet (liu2021spherical, ), PaiNN (schutt2021equivariant, ), PamNet (zhang2023universal, ) and LEFTNet (du2024new, ) etc.
@@
Due to that SchNet (schutt2017schnet, ), DimeNet++ (gasteiger_dimenetpp_2020, ), SphereNet (liu2021spherical, ), PaiNN (schutt2021equivariant, ), PamNet (zhang2023universal, ) and LEFTNet (du2024new, ) are trained under different train/val/test partitions, we rerun these six baselines, and report the results for a fair comparison. Additionally, we also deploy GMN  (huang2021equivariant, ) and SEGNO  (liu2023improving, ) on the molecule property prediction task and rerun EGNN (satorras2021n, ) and SEGNN (brandstetter2021geometric, ) under the same experimental settings. All these baselines are running on NVIDIA A100 GPUs. The experimental settings are as follows: Adam optimizer with learning rate ${1\times 10^{-3}}$, the number of epochs 1,000, input embedding dim 128, hidden embedding dim 128, output input embedding dim 128, batch size 128, weight decay ${1\times 10^{-12}}$, and layer number 7.
@@
We build upon the experimental setting introduced in (brandstetter2021geometric, ), where we aim to predict a variety of chemical properties for each of the molecules through regression. We optimized and report the Mean Absolute Error (MAE) between predictions and ground truth. In this section, we employ EGNN and SEGNN as the backbone of DuSEGO, named DuSEGO-EGNN and DuSEGO-SEGNN. In particular, EGNN and SEGNN are implemented based on the implementation details provided in (satorras2021n, ; brandstetter2021geometric, ), which are used as the coupling function for message passing (rusch2022graph, ). Then, we use the ODE solvers provided in torchdiffeq library for solving Eq 8 and Eq. 9.
@@
Table 5 depicts the overall experimental results of all methods on the QM9 dataset. Compared to the baseline methods, DuSEGO-EGNN achieves the best results for three energy variables ($\Delta\epsilon$, $\epsilon_{\text{HOMO}}$, and $\epsilon_{\text{LUMO}}$), while DuSEGO-SEGNN achieves the best results for four energy variables ($G$, $H$, $U$, and $U_{0}$). Additionally, we observe that incorporating second-order information into EGNN and SEGNN allows DuSEGO-EGNN and DuSEGO-SEGNN to perform better than before in most cases, outperforming GMN and SEGNO in all tasks, which highlights the effectiveness of our approach. As a result, DuSEGO-EGNN and DuSEGO-SEGNN generate competitive results in most property prediction tasks while maintaining relatively simple model structures, without incorporating additional data such as angles and spherical harmonics, as used in PAMNet, SphereNet, PaiNN, and LEFTNet.
@@ 7. Conclusion
In this paper, we introduced DuSEGO, a dual second-order equivariant graph neural network designed to address both feature and coordinate representations simultaneously through a graph ODE framework. Our approach not only maintains equivariant properties but also alleviates critical issues in deep graph neural networks, such as over-smoothing and gradient problems. Empirical results on dynamic system modeling, graph autoencoding, and molecular data tasks demonstrate DuSEGO’s superior performance over existing methods, especially in data efficiency and stability across different graph depths and sample sizes. These results underscore DuSEGO’s capacity to model complex physical and molecular systems effectively, even in limited-data scenarios. Future work could explore extending DuSEGO’s application to other domains that benefit from high-order motion laws or developing variations that address specific challenges in large-scale or sparse datasets. Additionally, further integration with third-order or higher-order dynamics could reveal more insights into complex graph-structured data, enhancing its versatility across a wider range of tasks.

### v3

@@ 6.3. Molecular Data
Finally, we evaluate DuSEGO against state-of-the-art baseline methods in molecular data, aiming to predict molecular chemical properties on 3D molecular graphs.
@@
The QM9 dataset is widely used in computational chemistry, which is specifically designed for the development and testing of quantum chemistry methods. It consists of approximately 134,000 stable small organic molecules made up of C, H, O, N, and F. Each molecule in the dataset contains up to nine non-hydrogen atoms. The dataset includes a diverse set of molecular properties that have been pre-computed using high-level quantum mechanical theories. We leveraged the QM9 dataset partitions from (anderson2019cormorant, ) and sampled 100,000 molecules for training, 18,000 for validation, and 13,000 for testing (brandstetter2021geometric, ). Additionally, all properties have been normalized by subtracting the mean and dividing by the Mean Absolute Deviation (satorras2021n, ).
@@
Various widely used methods for molecule property prediction are deployed for comparison, such as DimeNet++ (gasteiger_dimenetpp_2020, ), SphereNet (liu2021spherical, ), PaiNN (schutt2021equivariant, ), PamNet (zhang2023universal, ), and LEFTNet (du2024new, ). Additionally, we deploy GMN  (huang2021equivariant, ) and SEGNO  (liu2024segno, ) on the molecule property prediction task.
@@
Due to that DimeNet++ (gasteiger_dimenetpp_2020, ), SphereNet (liu2021spherical, ), PaiNN (schutt2021equivariant, ), PamNet (zhang2023universal, ), and LEFTNet (du2024new, ) are trained under different train/val/test partitions, we rerun these six baselines, and report the results for a fair comparison. We rerun EGNN (satorras2021n, ) and SEGNN (brandstetter2021geometric, ) under the same experimental settings. All these baselines are evaluated on NVIDIA A100 GPUs. The experimental settings are as follows: Adam optimizer with learning rate $1\times 10^{-3}$}, the number of epochs 1,000, input embedding dim 128, hidden embedding dim 128, output embedding dim 128, batch size 128, weight decay $1\times 10^{-12}$}, and layer number 7.
@@
We build upon the experimental setting introduced in (brandstetter2021geometric, ; satorras2021n, ), where we aim to predict a variety of chemical properties for each of the molecules through regression. We optimized and reported the Mean Absolute Error (MAE) between predictions and ground truth. In this section, we employ EGNN and SEGNN as the backbone of the DuSEGO method, named DuSEGO-EGNN and DuSEGO-SEGNN, respectively. In particular, EGNN and SEGNN are implemented based on the implementation details provided in (satorras2021n, ; brandstetter2021geometric, ), which are used as the coupling function for message passing (rusch2022graph, ). Again, we use the ODE solvers provided in the torchdiffeq library for solving Eq. 8 and Eq. 9.
@@
We use Table 6 to show the overall experimental results of all methods on the QM9 dataset. Compared to the baseline methods, our DuSEGO-EGNN achieves the best results for three energy variables ($\Delta\epsilon$, $\epsilon_{\text{HOMO}}$, and $\epsilon_{\text{LUMO}}$), while DuSEGO-SEGNN achieves the best results for four energy variables ($G$, $H$, $U$, and $U_{0}$). Additionally, we observe that incorporating second-order information into EGNN and SEGNN enables DuSEGO-EGNN and DuSEGO-SEGNN to outperform their original versions in most cases. In addition, DuSEGO-EGNN and DuSEGO-SEGNN outperform GMN and SEGNO in all tasks, which highlights the effectiveness of our approach. Therefore, we can conclude that DuSEGO-EGNN and DuSEGO-SEGNN can generate competitive results in most property prediction tasks while maintaining relatively simple model structures. To further clarify the conceptual distinction between our DuSEGO and specialized 3D architectures, we note that models such as PAMNet, SphereNet, PaiNN, and LEFTNet typically enhance expressiveness by explicitly incorporating higher-order geometric features or advanced attention mechanisms into their message-passing frameworks. In contrast, our ODE-based approach adopts a fundamentally different philosophy: DuSEGO models the evolution of node features and coordinates as a continuous-time dynamic process governed by second-order ODEs, providing a principled, physics-inspired mechanism for capturing complex geometric dependencies without requiring manual engineering of geometric priors. Notably, these two paradigms are not mutually exclusive. Domain-specific geometric features or attention modules can be integrated into the ODE-based framework to enhance its expressiveness further. However, as demonstrated by the results in Table 6, our approach achieves highly competitive performance across a broad range of property prediction tasks even without explicit geometric augmentation. This underscores the generality and simplicity of ODE-based dynamics, as well as their potential to complement specialized 3D architectures.
@@ 7. Conclusion
In this paper, we introduced DuSEGO, a dual second-order equivariant graph neural network designed to address both feature and coordinate representations simultaneously through a graph ODE framework. Our approach not only maintains equivariant properties but also addresses two critical issues of deep graph neural networks: over-smoothing and gradient explosion problems. Empirical results on dynamic system modeling, graph auto-encoding, and molecular data tasks demonstrate the superior performance of our proposed DuSEGO over existing methods, especially in terms of data efficiency and stability across different layer numbers of GNN. These results underscore DuSEGO’s capacity to model complex physical and molecular systems effectively. Future work could explore extending DuSEGO’s application to other domains that benefit from high-order motion laws or developing variations that address specific challenges in large-scale or sparse datasets. Additionally, further integration with higher-order dynamics could reveal more insights into complex graph-structured data, enhancing its versatility across a broader range of tasks.